The correct order of increasing C-O bond length of
CO, 〖CO〗_3^(2-), CO2 is
a. 〖CO〗_3^(2-) < CO2 < CO
b. CO2 < 〖CO〗_3^(2-) < CO
c. CO < 〖CO〗_3^(2-) < CO2
d. CO < CO2 < 〖CO〗_3^(2-)

Grade:11

3 Answers

Aditi Chauhan
9 years ago
Sol. Bond-length ∝ 1/(Bond order) Bond-order : CO2 = 2, CO = 3, 〖CO〗_3^(2-) = 1 + 1/3 = 4/3 Therefore, order of bond-length is 〖CO〗_3^(2-) < CO2 < CO
Roshan Dhadiwal
5 years ago
Bond length inversely proportional to bond order
B.O of CO=3,B.O of CO2=2,B.O of CO3^2-=4/3
Therefore order must be CO3^2->CO2>CO
D opt
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Bond-length ∝ 1/(Bond order)
Bond-order : CO2 = 2, CO = 3, 〖CO〗_3^(2-)
= 1 + 1/3
= 4/3
Therefore, order of bond-length is 〖CO〗_3^(2-) < CO2 < CO

Thanks and Regards